OpenStudy (anonymous):

find f^-1(x) f(x) = ln(x+1)

OpenStudy (jdoe0001):

well, if you change f(x) = ln(x+1) to exponent notation, you'd get \(\bf f(x) = ln(x+1)\implies y = log_e(x+1)\implies e^y=x+1\) now to find the inverse function, do like you'd with any other function, swap the variables and solve for "y"